# A Uniqueness Theorem for Distributed Computation under Physical Constraints

### v1

@@ Abstract.
$c$s: Computer systems organization Distributed architecturesccs: Theory of computation Models of computation
1. Introduction
Foundational models of computation, from the Turing Machine and Lambda Calculus to concurrent paradigms like the Actor Model(Agha, 1986) and Process Calculi, have provided a powerful and abstract lens through which to understand the logical limits and expressive power of computation. A common thread uniting these models is their deliberate abstraction from the physical realities of the underlying hardware; they often assume, for instance, unbounded memory, flawless communication channels, or negligible energy costs. This paper investigates a fundamental question that arises when this abstraction is removed: What essential form must a distributed computational paradigm take when it is forced to operate under a set of inviolable physical axioms, such as strict memory bounds and communication lower bounds, that are inherent to its execution environment? This question is no longer a purely theoretical curiosity but a pressing challenge, brought into sharp focus by the emergence of highly constrained, massively parallel computing environments such as in-network computing.
Nowhere are these physical constraints more acute than in the domain of In-Network Computing (INC). The advent of programmable network devices, such as P4 switches and Data Processing Units (DPUs), has presented an unprecedented opportunity for line-rate data processing, but this new paradigm operates under conditions that fundamentally distinguish it from traditional distributed computing. Specifically, a devices’s fast memory is severely limited and physically independent of the total problem size ($W_{p}=O(1)$), while the network environment it inhabits is inherently unreliable. These physical laws impose a stringent, unavoidable trilemma: the concurrent optimization of communication efficiency, bounded memory, and robust scalability. Prevailing paradigms like the Bulk Synchronous Parallel (BSP) model (Valiant, 1990) or stateful message-passing, designed for more resource-rich environments, falter under these conditions, which necessitates a return to first principles to identify the viable computational models that can operate within such severe limitations.
@@ 1. Introduction
2. Theoretical Context and Related Models of Computation
@@
2.1. In-Network Computing Models and Systems
The field of In-Network Computing has produced several powerful programming models and systems, such as (Bosshart et al., 2014) and various frameworks for DPUs. These systems provide essential abstractions for directly programming the network data plane. However, their primary focus is often on expressing packet processing logic and achieving line-rate performance. Issues like robust scalability, fault tolerance, and memory management under heavy traffic are typically addressed through implementation-specific, ad-hoc techniques rather than a guiding theoretical framework.
These systems, therefore, serve as a compelling manifestation of the problem we address: they are powerful engineering artifacts whose necessary design patterns and trade-offs have been discovered through practice, but without a formal theory to explain their inevitability. This paper provides that missing theoretical foundation, deriving from first principles the unique paradigm (SDPF) that must emerge under the physical laws governing these systems.(Kianpisheh and Taleb, 2022; Hauser et al., 2023; Bosshart et al., 2014)
2.2. Communication and I/O Lower Bounds
In the INC context, the on-chip memory of a network device and the hosts main memory form a classic two-level storage hierarchy, making communication lower bounds acutely relevant. Classic works, using models such as the red-blue pebble game and communication-avoiding algorithms, have characterized the unavoidable communication costs in such systems (e.g., (Jia-Wei and Kung, 1981; Irony et al., 2004; Ballard et al., 2013)).
@@ 2.2. Communication and I/O Lower Bounds
2.3. Logical Convergence: CALM and Semilattices
The challenge of achieving consistency without costly coordination has been masterfully addressed at the logical level. The CALM (Consistency as Logical Monotonicity) theorem provides the fundamental principle: monotonic programs are guaranteed to be eventually consistent without coordination (Alvaro et al., 2011). This principle finds its algebraic manifestation in semilattice structures, as elegantly formalized by Conflict-free Replicated Data Types (CRDTs). CRDTs leverage commutative, associative, and idempotent merge operators to ensure convergence regardless of network delays, reordering, or duplication (Shapiro et al., 2011).
@@ 2.3. Logical Convergence: CALM and Semilattices
2.4. Operational Models and their Physical Infeasibility
@@
While the models discussed above address logical convergence, the dominant operational paradigms in distributed computing were architected with different, more forgiving physical assumptions. Paradigms like Valiants Bulk Synchronous Parallel (BSP) model, and its descendants in batch (Dean and Ghemawat, 2008; Zaharia et al., 2012) and streaming (Flink/Beam), rely on coordination mechanisms such as global barriers, checkpoints, or ordered watermarks (Akidau et al., 2015; Lamport, 2019). As formally proven in our Proposition 5.3, any form of global synchronization is fundamentally incompatible with Axiom A2 (Perturbation and Failures), as heavy-tailed latency inevitably destroys linear scalability. These paradigms are therefore physically infeasible in the INC domain.
@@
Similarly, classic logical models for concurrency, while foundational, abstract away the very constraints we consider paramount. Kahn Process Networks (KPN) guarantee determinism through the assumption of unbounded FIFO channels, a direct violation of our Axiom A3 (Memory Upper Bound) (Gilles, 1974). The Actor model, a powerful abstraction for composable concurrency, focuses on behavioral guarantees and does not inherently address the physical lower bounds of communication (Axiom A1) or memory (Axiom A3) (Agha, 1986). Our work does not refute the logical elegance of these models; instead, by elevating physical laws to first-class axioms, we prove the necessary operational form a concurrent model must take when it can no longer ignore these harsh realities.
@@
More advanced dataflow systems, such as Naiad (Murray et al., 2013) or Differential Dataflow (McSherry et al., 2013), have replaced coarse-grained barriers with more sophisticated progress tracking mechanisms like timestamps or epochs. While this represents a significant improvement, these mechanisms still constitute a form of logical coordination that is fundamentally vulnerable within our axiomatic regime. A single straggler suffering from heavy-tailed latency (Axiom A2) can halt the progress of a timestamp or epoch, preventing the system from reclaiming resources or committing results. This reintroduces a subtle form of the convoy effect and can lead to memory pressure that violates the strict bounded memory guarantee (Axiom A3), as the system is forced to buffer ever-increasing amounts of out-of-order data. Thus, even these state-of-the-art models are incompatible with the strict physical realities we study.
2.5. A Constructive Duality to Impossibility Theorems
@@
Classic impossibility results, most notably the FLP and CAP theorems, define the fundamental boundaries of what is achievable in asynchronous distributed systems facing failures (Fischer et al., 1985; Gilbert and Lynch, 2002). Our work offers a constructive dual: rather than proving what is impossible to achieve (e.g., consensus), we prove what is inevitable for a different class of problems. We start by accepting a set of axioms, including one that explicitly circumvents the need for consensus (Axiom A4), and derive the unique computational form that is not just possible, but necessary.
@@
This perspective provides a formal underpinning for navigating challenges like the tail at scale phenomenon (Dean and Barroso, 2013). In the physically-constrained environments we study, a single straggler, a direct consequence of the heavy-tailed perturbations formalized in Axiom A2, is not merely a performance issue but a threat to system survival due to bounded memory (Axiom A3). While impossibility theorems erect walls, our uniqueness theorem illuminates the single, narrow path that remains. By proving from first principles that barrier-less, data-centric computation is a condition for survival, we provide a constructive guide for building robust systems in this challenging domain.
2.6. Summary of Theoretical Positioning
@@
Table 1. Comparison of Computational Models Against Core Axiomatic Requirements
@@
This Works Goal Yes Yes Yes Yes
BSP Model (Valiant, 1990) No No No No
KPN (Gilles, 1974) Yes No Yes
Actor Model (Agha, 1986) Yes Yes
Dataflow (Murray et al., 2013) No No No No
CRDTs (Shapiro et al., 2011) Yes Yes
@@
3. Formal Model and Axioms
@@
3.1. Execution Model and Cost Metrics
@@ 3.2. Axiomatic System
4. The SDPF Paradigm
(a) Process-Centric Model
(b) The SDPF Data-Centric Model
Figure 1. A conceptual comparison between traditional process-centric models and the data-centric SDPF paradigm. In SDPF, intelligence and control logic are embedded in the data flow, while executors are stateless and anonymous.
@@ 4. The SDPF Paradigm
4.1. Single-Read Reuse (S1)
@@
4.2. Stateless Micro-tasks (S2)
@@ 4.3. Idempotent Merge (S3)
4.4. Barrier-less Asynchronous Scheduling (S4)
@@ 4.5. Sliding Window (S5)
5. Proof of Necessity and Uniqueness
@@
Table 2. Proof Roadmap for SDPF Uniqueness
@@
Communication Efficiency Single-Read Reuse I/O is amplified A1 (Comm. Lower Bound) Proposition 5.1
Robustness Stateless/Idempotent Fault tolerance is costly A1 / A2 Proposition 5.2
Scalability Barrier-less Async. Tail latency is amplified A2 (Perturbation Model) Proposition 5.3
Bounded Memory Sliding Window Memory is coupled with scale A3 (Memory Upper Bound) Proposition 5.4
@@
5.1. Necessary Condition for Communication Efficiency
@@
Proof Sketch.
This proof aims to quantify the communication overhead incurred by violating the single-read reuse principle. Our strategy is to express the lower bound of the total communication volume $Q$ as the sum of two components: the unavoidable baseline I/O, which is $|I|+|O|$, and the redundant I/O generated by each tile $\text{tile}_{k}$ being repeatedly fetched from the source $U_{k}-1$ times. By substituting this lower bound of $Q$ into the definition of the I/O amplification factor $R_{A}$, we can directly derive that $R_{A}$ must be strictly greater than 1. The excess part is proportional to the normalized redundant communication volume. To satisfy the communication optimality Axiom 1 (i.e., $R_{A}=O(1)$), this redundant term must be eliminated, thus proving the necessity of single-read reuse (S1). The detailed formal proof is provided in Appendix A. ∎
5.2. Necessary Conditions for Robustness
@@
Proof Sketch.
This proof aims to reveal that, in the environment of perturbations and failures described by Axiom 2, any paradigm relying on mutable state or non-idempotent operations inevitably falls into a dilemma. To ensure correctness, such systems must choose between two costly fault-tolerance strategies. The first is to remember through persistence (e.g., checkpoints), which generates enormous communication overhead, violating Axiom 1. The second is to enforce order (e.g., total-order broadcast), which is equivalent to a global barrier and, according to Proposition 5.3, destroys linear scalability. Since both paths lead to the violation of core axioms, the paradigm must adopt stateless micro-tasks (S2) and idempotent merge (S3). The detailed formal proof is provided in Appendix B. ∎
@@
Proof Sketch.
This proof reveals from first principles why synchronization barriers combined with heavy-tailed latency (Axiom 2) fundamentally destroy linear scalability. The convoy effect of barriers means the round time is determined by the slowest task. Based on principles from extreme value theory, for any heavy-tailed distribution, the expectation of the maximum of $|P|$ i.i.d. variables is a monotonically increasing function of $|P|$. This is in stark contrast to distributions with exponentially decaying tails (e.g., Gaussian), where the expected maximum grows much more slowly, typically logarithmically. Since throughput $S(P)\propto|P|/E[T_{\mathrm{round}}]$, an increasing round time necessarily leads to sub-linear scalability ($S(P)=o(P)$). Therefore, barrier-less scheduling (S4) is a necessary condition. The detailed formal proof is provided in Appendix C. ∎
5.3. Necessary Condition for Bounded Memory
@@
Proof Sketch.
This proof originates from I/O complexity theory, which reveals a fundamental trade-off between memory size and communication volume. To approach the communication lower bound (Axiom 1), an algorithm must maximize data reuse, which incentivizes hoarding data in memory. However, I/O complexity theory proves that for many problems, the size of the live dataset required to achieve optimal communication grows polynomially with the problem size $n$. This leads to $W_{\max}=\Omega(n^{\beta})$, which directly contradicts the bounded memory Axiom 3 ($W_{\max}=O(1)$). The only way to resolve this conflict is to adopt a use and discard pipelined model, which is the essence of the sliding window (S5). The detailed formal proof is provided in Appendix D. ∎
5.4. The Uniqueness Main Theorem
@@
Proof Sketch.
This proof employs a constructive reduction method. We start with an arbitrary optimal paradigm $\mathcal{E}$ and apply a series of transformations (T1-T4). Propositions 5.1 through 5.4 provide the theoretical basis, showing that each correction is necessary to satisfy a core axiom. Crucially, each transformation is shown to be performance-preserving. The overheads introduced, such as the management of replay identifiers (rid) in T2 or the buffering required by T4, are shown to be constant or bounded, and most importantly, independent of the system scale $|P|$ and network topology. The resulting paradigm $\mathcal{E}$ is, by definition, an SDPF normal form and is performance-equivalent to $\mathcal{E}$. This proves that SDPF is the unique normal form for all optimal paradigms. The full proof with formal preservation lemmas for each transformation is provided in Appendix E. ∎
Figure 2. The constructive reduction proof of Theorem $5.$ Any paradigm claiming optimality under the axioms is progressively transformed by eliminating inefficiencies (T1-T4), inevitably converging to the unique SDPF normal form.
@@
6. Convergence, Completeness, and Minimality

### v2

# Uniqueness Theorem for Distributed Computation

@@ Abstract
1 Introduction
Foundational models of computation, from the Turing Machine and Lambda Calculus to concurrent paradigms like the Actor Model[1] and Process Calculi, have provided a powerful and abstract lens through which to understand the logical limits and expressive power of computation. A common thread uniting these models is their deliberate abstraction from the physical realities of the underlying hardware; they often assume, for instance, unbounded memory, flawless communication channels, or negligible energy costs. This paper investigates a fundamental question that arises when this abstraction is removed: What essential form must a distributed computational paradigm take when it is forced to operate under a set of inviolable physical axioms, such as strict memory bounds and communication lower bounds, that are inherent to its execution environment? This question is no longer a purely theoretical curiosity but a pressing challenge, brought into sharp focus by the emergence of highly constrained, massively parallel computing environments such as in-network computing.
Nowhere are these physical constraints more acute than in the domain of In-Network Computing (INC). The advent of programmable network devices, such as P4 switches and Data Processing Units (DPUs), has presented an unprecedented opportunity for line-rate data processing, but this new paradigm operates under conditions that fundamentally distinguish it from traditional distributed computing. Specifically, a devices’s fast memory is severely limited and physically independent of the total problem size ($W_{p}=O(1)$), while the network environment it inhabits is inherently unreliable. These physical laws impose a stringent, unavoidable trilemma: the concurrent optimization of communication efficiency, bounded memory, and robust scalability. Prevailing paradigms like the Bulk Synchronous Parallel (BSP) model [2] or stateful message-passing, designed for more resource-rich environments, falter under these conditions, which necessitates a return to first principles to identify the viable computational models that can operate within such severe limitations.
@@ 1 Introduction
2 Theoretical Context and Related Models of Computation
@@
2.1 In-Network Computing Models and Systems
The field of In-Network Computing has produced several powerful programming models and systems, such as [3] and various frameworks for DPUs. These systems provide essential abstractions for directly programming the network data plane. However, their primary focus is often on expressing packet processing logic and achieving line-rate performance. Issues like robust scalability, fault tolerance, and memory management under heavy traffic are typically addressed through implementation-specific, ad-hoc techniques rather than a guiding theoretical framework.
These systems, therefore, serve as a compelling manifestation of the problem we address: they are powerful engineering artifacts whose necessary design patterns and trade-offs have been discovered through practice, but without a formal theory to explain their inevitability. This paper provides that missing theoretical foundation, deriving from first principles the unique paradigm (SDPF) that must emerge under the physical laws governing these systems.[4, 5, 3]
2.2 Communication and I/O Lower Bounds
In the INC context, the on-chip memory of a network device and the hosts main memory form a classic two-level storage hierarchy, making communication lower bounds acutely relevant. Classic works, using models such as the red-blue pebble game and communication-avoiding algorithms, have characterized the unavoidable communication costs in such systems (e.g., [6, 7, 8]).
@@ 2.2 Communication and I/O Lower Bounds
2.3 Logical Convergence: CALM and Semilattices
The challenge of achieving consistency without costly coordination has been masterfully addressed at the logical level. The CALM (Consistency as Logical Monotonicity) theorem provides the fundamental principle: monotonic programs are guaranteed to be eventually consistent without coordination [9]. This principle finds its algebraic manifestation in semilattice structures, as elegantly formalized by Conflict-free Replicated Data Types (CRDTs). CRDTs leverage commutative, associative, and idempotent merge operators to ensure convergence regardless of network delays, reordering, or duplication [10].
@@ 2.3 Logical Convergence: CALM and Semilattices
2.4 Operational Models and their Physical Infeasibility
@@
While the models discussed above address logical convergence, the dominant operational paradigms in distributed computing were architected with different, more forgiving physical assumptions. Paradigms like Valiants Bulk Synchronous Parallel (BSP) model, and its descendants in batch [11, 12] and streaming (Flink/Beam), rely on coordination mechanisms such as global barriers, checkpoints, or ordered watermarks [13, 14]. As formally proven in our Proposition 5.3, any form of global synchronization is fundamentally incompatible with Axiom A2 (Perturbation and Failures), as heavy-tailed latency inevitably destroys linear scalability. These paradigms are therefore physically infeasible in the INC domain.
@@
Similarly, classic logical models for concurrency, while foundational, abstract away the very constraints we consider paramount. Kahn Process Networks (KPN) guarantee determinism through the assumption of unbounded FIFO channels, a direct violation of our Axiom A3 (Memory Upper Bound) [15]. The Actor model, a powerful abstraction for composable concurrency, focuses on behavioral guarantees and does not inherently address the physical lower bounds of communication (Axiom A1) or memory (Axiom A3) [1]. Our work does not refute the logical elegance of these models; instead, by elevating physical laws to first-class axioms, we prove the necessary operational form a concurrent model must take when it can no longer ignore these harsh realities.
@@
More advanced dataflow systems, such as Naiad [16] or Differential Dataflow [17], have replaced coarse-grained barriers with more sophisticated progress tracking mechanisms like timestamps or epochs. While this represents a significant improvement, these mechanisms still constitute a form of logical coordination that is fundamentally vulnerable within our axiomatic regime. A single straggler suffering from heavy-tailed latency (Axiom A2) can halt the progress of a timestamp or epoch, preventing the system from reclaiming resources or committing results. This reintroduces a subtle form of the convoy effect and can lead to memory pressure that violates the strict bounded memory guarantee (Axiom A3), as the system is forced to buffer ever-increasing amounts of out-of-order data. Thus, even these state-of-the-art models are incompatible with the strict physical realities we study.
2.5 A Constructive Duality to Impossibility Theorems
@@
Classic impossibility results, most notably the FLP and CAP theorems, define the fundamental boundaries of what is achievable in asynchronous distributed systems facing failures [18, 19]. Our work offers a constructive dual: rather than proving what is impossible to achieve (e.g., consensus), we prove what is inevitable for a different class of problems. We start by accepting a set of axioms, including one that explicitly circumvents the need for consensus (Axiom A4), and derive the unique computational form that is not just possible, but necessary.
@@
This perspective provides a formal underpinning for navigating challenges like the tail at scale phenomenon [20]. In the physically-constrained environments we study, a single straggler, a direct consequence of the heavy-tailed perturbations formalized in Axiom A2, is not merely a performance issue but a threat to system survival due to bounded memory (Axiom A3). While impossibility theorems erect walls, our uniqueness theorem illuminates the single, narrow path that remains. By proving from first principles that barrier-less, data-centric computation is a condition for survival, we provide a constructive guide for building robust systems in this challenging domain.
2.6 Summary of Theoretical Positioning
@@
Table 1: Comparison of Computational Models Against Core Axiomatic Requirements
@@
  This Work’s Goal   Yes   Yes   Yes   Yes
  BSP Model [2]   No   No   No   No
  KPN [15]   –   Yes   No   Yes
  Actor Model [1]   –   Yes   –   Yes
  Dataflow [16]   No   No   No   No
  CRDTs [10]   –   Yes   –   Yes
@@
3 Formal Model and Axioms
@@
3.1 Execution Model and Cost Metrics
@@ 3.2 Axiomatic System
4 The SDPF Paradigm
Figure 1: A conceptual comparison between traditional process-centric models and the data-centric SDPF paradigm. (a) Process-Centric Model; (b) The SDPF Data-Centric Model. In SDPF, intelligence and control logic are embedded in the data flow, while executors are stateless and anonymous.
@@ 4 The SDPF Paradigm
4.1 Single-Read Reuse (S1)
@@
4.2 Stateless Micro-tasks (S2)
@@ 4.3 Idempotent Merge (S3)
4.4 Barrier-less Asynchronous Scheduling (S4)
@@ 4.5 Sliding Window (S5)
5 Proof of Necessity and Uniqueness
@@
Table 2: Proof Roadmap for SDPF Uniqueness
@@
Communication Efficiency Single-Read Reuse I/O is amplified A1 (Comm. Lower Bound) Prop. 5.1
Robustness Stateless/Idempotent Fault tolerance is costly A1 / A2 Prop. 5.2
Scalability Barrier-less Async. Tail latency is amplified A2 (Perturbation Model) Prop. 5.3
Bounded Memory Sliding Window Memory is coupled with scale A3 (Memory Upper Bound) Prop. 5.4
@@
5.1 Necessary Condition for Communication Efficiency
@@
Proof 5.2 (Proof Sketch).
This proof aims to quantify the communication overhead incurred by violating the single-read reuse principle. Our strategy is to express the lower bound of the total communication volume $Q$ as the sum of two components: the unavoidable baseline I/O, which is $|I|+|O|$, and the redundant I/O generated by each tile $\text{tile}_{k}$ being repeatedly fetched from the source $U_{k}-1$ times. By substituting this lower bound of $Q$ into the definition of the I/O amplification factor $R_{A}$, we can directly derive that $R_{A}$ must be strictly greater than 1. The excess part is proportional to the normalized redundant communication volume. To satisfy the communication optimality Axiom 1 (i.e., $R_{A}=O(1)$), this redundant term must be eliminated, thus proving the necessity of single-read reuse (S1). The detailed formal proof is provided in A.
5.2 Necessary Conditions for Robustness
@@
Proof 5.4 (Proof Sketch).
This proof aims to reveal that, in the environment of perturbations and failures described by Axiom 2, any paradigm relying on mutable state or non-idempotent operations inevitably falls into a dilemma. To ensure correctness, such systems must choose between two costly fault-tolerance strategies. The first is to remember through persistence (e.g., checkpoints), which generates enormous communication overhead, violating Axiom 1. The second is to enforce order (e.g., total-order broadcast), which is equivalent to a global barrier and, according to Proposition 5.5, destroys linear scalability. Since both paths lead to the violation of core axioms, the paradigm must adopt stateless micro-tasks (S2) and idempotent merge (S3). The detailed formal proof is provided in B.
@@
Proof 5.6 (Proof Sketch).
This proof reveals from first principles why synchronization barriers combined with heavy-tailed latency (Axiom 2) fundamentally destroy linear scalability. The convoy effect of barriers means the round time is determined by the slowest task. Based on principles from extreme value theory, for any heavy-tailed distribution, the expectation of the maximum of $|P|$ i.i.d. variables is a monotonically increasing function of $|P|$. This is in stark contrast to distributions with exponentially decaying tails (e.g., Gaussian), where the expected maximum grows much more slowly, typically logarithmically. Since throughput $S(P)\propto|P|/E[T_{\mathrm{round}}]$, an increasing round time necessarily leads to sub-linear scalability ($S(P)=o(P)$). Therefore, barrier-less scheduling (S4) is a necessary condition. The detailed formal proof is provided in C.
5.3 Necessary Condition for Bounded Memory
@@
Proof 5.8 (Proof Sketch).
This proof originates from I/O complexity theory, which reveals a fundamental trade-off between memory size and communication volume. To approach the communication lower bound (Axiom 1), an algorithm must maximize data reuse, which incentivizes hoarding data in memory. However, I/O complexity theory proves that for many problems, the size of the live dataset required to achieve optimal communication grows polynomially with the problem size $n$. This leads to $W_{\max}=\Omega(n^{\beta})$, which directly contradicts the bounded memory Axiom 3 ($W_{\max}=O(1)$). The only way to resolve this conflict is to adopt a use and discard pipelined model, which is the essence of the sliding window (S5). The detailed formal proof is provided in D.
5.4 The Uniqueness Main Theorem
@@
Proof 5.10 (Proof Sketch).
This proof employs a constructive reduction method. We start with an arbitrary optimal paradigm $\mathcal{E}$ and apply a series of transformations (T1-T4). Propositions 5.1 through 5.7 provide the theoretical basis, showing that each correction is necessary to satisfy a core axiom. Crucially, each transformation is shown to be performance-preserving. The overheads introduced, such as the management of replay identifiers (rid) in T2 or the buffering required by T4, are shown to be constant or bounded, and most importantly, independent of the system scale $|P|$ and network topology. The resulting paradigm $\mathcal{E}$ is, by definition, an SDPF normal form and is performance-equivalent to $\mathcal{E}$. This proves that SDPF is the unique normal form for all optimal paradigms. The full proof with formal preservation lemmas for each transformation is provided in E.
Figure 2: The constructive reduction proof of Theorem $5.$ Any paradigm claiming optimality under the axioms is progressively transformed by eliminating inefficiencies (T1-T4), inevitably converging to the unique SDPF normal form.
@@
6 Convergence, Completeness, and Minimality